# The Unbounded Denominators Conjecture

### v3

@@ 4.3. A leveraging argument
Let us assume that there exists an (even) $N$ such that $R_{N}$ is strictly larger than $M_{N}$. Let $f(\tau)\in\mathbf{Z}\llbracket q^{1/N}\rrbracket\in R_{N}$ be an element which does not lie in $M_{N}$. Recall that all forms in $R_{N}$ and thus in particular $f$ is invariant by a subgroup $G=G_{N}\subset\langle E,\Gamma(N)\rangle$ which is normal with finite index in $\langle E,\Gamma(N)\rangle$ and has $L(G_{N})=N$ by Lemma 4.2.3. The main idea of this section is to exploit the fact that $f(p\tau)\in\mathbf{Z}\llbracket q^{1/N}\rrbracket$ is also a modular form with integer coefficients for any prime $p$. Since the form $f(\tau)$ is invariant under $G$, the form $f(p\tau)$ is invariant under $A^{-1}GA$ and thus also the group $A^{-1}GA\cap\mathrm{SL}_{2}(\mathbf{Z})$. Now, by Lemma 4.1.7, we know that this group has (Wohlfahrt) level dividing $Np$. In particular $f(p\tau)$ has cusp width dividing $Np$ at each cusp, and hence $f(p\tau)\in R_{Np}$.
@@
Suppose that $(p,N)=1$ is prime. Suppose that $f(\tau)$ is not invariant under a congruence subgroup. Then the form $f(p\tau)$ is not in the $M_{Np}$-algebra generated by $R_{N}$.
@@
Suppose that there exists an $N$ such that $R_{N}:M_{N}]>1$. Then for any prime $p$ not dividing $N$, one has
$R_{Np}:M_{Np}]\geq 2[R_{N}:M_{N}].$
@@
Let $f(\tau)$ be a form in $R_{Np}$ which is not in $M_{N}$. By Theorem 4.3.1, we deduce that $f(p\tau)\in R_{Np}$ is not in the $M_{Np}$-algebra $R_{N}M_{Np}$ generated by $R_{N}$ (which is a subfield of $R_{Np}$). We have
$R_{Np}:M_{N}]=[R_{Np}:R_{N}M_{Np}][R_{N}M_{Np}:M_{N}]=[R_{Np}:R_{N}M_{Np}][R_{N}:M_{N}][M_{Np}:M_{N}],$
because the intersection of $M_{Np}$ and $R_{N}$ is $M_{N}$. Thus
$\frac{[R_{Np}:M_{Np}]}{[R_{N}:M_{N}]}=[R_{Np}:R_{N}M_{Np}]$
@@
Our goal is to prove that $R_{N}=M_{N}$. We can and do assume that $N$ is even. As noted in Lemma 4.2.3, $R_{N}:M_{2}]<\infty$. The degree of $M_{N}$ over $M_{2}$ is equal to the degree of the modular curve $Y(N)$ over $Y(2)$, and this is given by the explicit formula
(4.3.3) $=\frac{1}{2}[\Gamma(2):\Gamma(N)]=$ $\ \frac{N^{3}}{2[\mathrm{SL}_{2}(\mathbf{Z}):\Gamma(2)]}\prod_{p|N}\left(1-\frac{1}{p^{2}}\right)$
$>$ $\ \frac{N^{3}}{2[\mathrm{SL}_{2}(\mathbf{Z}):\Gamma(2)]}\prod_{p}\left(1-\frac{1}{p^{2}}\right)=\frac{N^{3}}{12\zeta(2)}.$
Since $R_{N}:M_{2}]=[R_{N}:M_{N}]\cdot[M_{N}:M_{2}]$, it follows that we have a bound:
(4.3.4) $R_{N}:M_{N}]\leq\frac{12\zeta(2)[R_{N}:M_{2}]}{N^{3}}$
for all $N$. We can now compare this bound against the one coming from Theorem 4.3.2
@@
$R_{N}:M_{2}]\leq CN^{3}\log N$
for all even integers $N$. Then $R_{N}=M_{N}$ for every $N$, that is, the unbounded denominators conjecture holds.
@@
Assume there exists an $N$ such that $R_{N}\neq M_{N}$. Let $S$ denote the set of primes <X<X which are co-prime to $N$. By induction, Theorem 4.3.2 implies for such an $N$ that
(4.3.6) $R_{N\prod_{p\in S}p}:M_{N\prod_{p\in S}p}]\geq 2^{\#S}>2^{(1-\varepsilon)X/\log X},$
by the prime number theorem. The RHS certainly increases faster than any power of $X$. On the other hand, from the assumed bound on $R_{N}:M_{2}]$ together with the bound (4.3.4), we obtain
(4.3.7) $[R_{N\prod_{p\in S}p}:M_{N\prod_{p\in S}p}]\leq$ $12C\zeta(2)\log\left(N\prod_{p\in S}p\right)$
[…]
where the last inequality follows from the prime number theorem. Combining the bounds (4.3.6) and (4.3.7) gives
$2^{(1-\varepsilon)X/\log X}<12C\zeta(2)X(1+\varepsilon),$
which (by some margin!) is a contradiction for sufficiently large $X$. ∎
@@
The argument still works with a bound weaker than $R_{N}:M_{2}]\ll N^{3}\log N$, although $R_{N}:M_{2}]\ll N^{3+\varepsilon}$ would not be strong enough.
@@ 4.4. Amalgams and a non-abelian version of Ihara’s Lemma
This is more or less precisely [Tho89, Theorem 3] with the addition of level structure as in [Ber94]. One may think of Lemma 4.4.1 as a non-abelian version of Ihara’s Lemma, because (as explained below in the proof of Lemma 4.6.2) the case when $S$ is a vector space over $\mathbf{F}_{q}$ reduces to the statement of Ihara’s Lemma as proved by Ribet [Rib84]. (The proofs of both claims are very similar.)
@@
The intersection of $\langle E,\Gamma(N)\rangle$ with $A\langle E,\Gamma(N)\rangle A^{-1}$ is the group $\langle E,\Gamma(N)\rangle\cap\Gamma_{0}(p)$. We proceed by contradiction. Assume that the map $(f_{1},f_{2})$ is not surjective. By Goursat’s lemma, there exists a non-trivial quotient $\Delta$ of $S$ and projections $\pi_{i}:S\rightarrow\Delta$ such that the composites $\pi_{1}\circ f_{1}$ and $\pi_{2}\circ f_{2}$ agree. We define a map $g_{1}$ by the composite
@@
Let $\gamma=\left(\begin{matrix}1&N\\ 0&1\end{matrix}\right)$ and $\eta=\left(\begin{matrix}1&0\\ N&1\end{matrix}\right)$. The assumption that $L(G)=N$ and $G$ has finite index in $\mathrm{SL}_{2}(\mathbf{Z})$ impies that $\gamma,\eta\in G$. Since $A^{-1}\gamma^{p}A^{-1}=\gamma\in A^{-1}GA\cap\Gamma_{0}(p)$, we see that $\gamma\in\ker(f_{1})$ and $\gamma\in\ker(f_{2})$, and yet
@@
(4.4.5) $H^{1}(\Gamma(N)/\mathrm{S}G,\mathbf{F}_{q})^{\oplus 2}=H^{1}(S,\mathbf{F}_{q})^{2}\rightarrow H^{1}(\Gamma(N)\cap\Gamma_{0}(p),\mathbf{F}_{q}),$
@@
where $H^{1,\mathrm{cong}}(\Gamma(N),\mathbf{F}_{q})\subset H^{1}(\Gamma(N),\mathbf{F}_{q})$ denotes the classes which vanish after restriction to a congruence subgroup (Definition 4.5.1). This is because the kernel of any non-trivial map in $H^{1,\mathrm{cong}}(\Gamma(N),\mathbf{F}_{q})$ has level strictly divisible by $N$. The claim (4.4.5) follows from (4.4.6) as a consequence of Ihara’s Lemma, as proved by Ribet [Rib84] (see Lemma 4.6.2). The maps $\langle E,\Gamma(N)\rangle\cap\Gamma_{0}(p)\rightarrow B/\langle U\rangle\rightarrow\mathbf{F}_{q}$ on the other hand come from the classes in $H^{1}(\Gamma(N)\cap\Gamma_{0}(p),\mathbf{F}_{q})$ which restricts to zero on $H^{1}(\Gamma(N)\cap\Gamma_{1}(p),\mathbf{F}_{q})$, and thus what is required is to upgrade the injection of (4.4.5) to an injection
(4.4.7) $H^{1}(\Gamma(N)/\mathrm{S}G,\mathbf{F}_{q})^{\oplus 2}=H^{1}(S,\mathbf{F}_{q})^{2}\rightarrow H^{1}(\Gamma(N)\cap\Gamma_{1}(p),\mathbf{F}_{q}),$
@@
Because $\gamma$ and $E$ map to zero in $S\times S$ and $B/\langle E,U\rangle$ has order (p1)/2(p-1)/2, The proof of Lemma 4.4.4 is almost immediate if one imposes the additional hypothesis that $\left(\frac{p-1}{2},|S|\right)=1$. In particular, one would not have to appeal to the results in §4.5 and §4.6 (which are not used elsewhere in this paper). It turns out that proving Lemma 4.4.4 under this weaker hypothesis would suffice for the proof of the unbounded denominators conjecture. The key point is that if $\langle E,\Gamma(N)\rangle/G_{N}\simeq S$ and $G_{Np}$ is the group given by the intersection of the three groups $\langle E,\Gamma(Np)\rangle$, $G$, and $AGA^{-1}$, then $\langle E,\Gamma(Np)\rangle/G_{Np}\simeq S\times S$. In particular, one can control the primes dividing $S$ as one varies $N$. Then, in the argument of Proposition 4.3.5, instead of adding all primes <X<X prime to $N$, one only includes primes in some arithmetic progression satisfying the congruence $\left(\frac{p-1}{2},|S|\right)=1$ for some fixed $S$. However, it seems more natural to prove Lemma 4.4.4 without such an ugly hypothesis. Additionally, sections §4.5 and §4.6 may be of independent interest.
Returning to the assumptions of Lemma 4.4.4, let $K=\langle\ker((f_{1},\pi),\ker(f_{2})\rangle$ be the group generated by $\ker((f_{1},\pi))$ and $\ker(f_{2})$. We deduce from Lemma 4.4.4 that the image of $(f_{1},f_{2},\pi)$ contains the elements $(x,0,z)$ and $(0,y,0)$ for any triple $(x,y,z)\in S\times S\times B$. But the pre-images of these elements clearly lie in $\ker(f_{2})$ and $\ker((f_{1},\pi))$ respectively, and thus lie in $K$. But then the pre-image of any element lies in $K$, and we deduce that $K=\langle E,\Gamma(N)\rangle\cap\Gamma_{0}(p)$, or equivalently that
@@
Consider the function $f(p\tau)$. Assume that this lies in the algebra generated by $f(\tau)$ and $M_{Np}$. Then $f(p\tau)$ is invariant under both $A^{-1}GA\cap\mathrm{SL}_{2}(\mathbf{Z})$ and $G\cap\Gamma(Np)$. But from (4.4.9) we see that these groups together generate a congruence subgroup, and thus $f(p\tau)$ and $f(\tau)$ are congruence, a contradiction. ∎

### v4

@@ 4.3. A leveraging argument
Let us assume that there exists an $N$ such that $R_{2N}$ is strictly larger than $M_{2N}$. Let $f(\tau)\in\mathbf{Z}\llbracket q^{1/N}\rrbracket\in R_{2N}$ be an element which does not lie in $M_{2N}$. Recall that all forms in $R_{2N}$ and thus in particular $f$ is invariant by a subgroup $G=G_{2N}\subset\langle E,\Gamma(2N)\rangle$ which is normal with finite index in $\langle E,\Gamma(2N)\rangle$ and has $L(G_{2N})=2N$ by Lemma 4.2.3. The main idea of this section is to exploit the fact that $f(p\tau)\in\mathbf{Z}\llbracket q^{1/N}\rrbracket$ is also a modular form with integer coefficients for any prime $p$. Since the form $f(\tau)$ is invariant under $G$, the form $f(p\tau)$ is invariant under $A^{-1}GA$ and thus also the group $A^{-1}GA\cap\mathrm{SL}_{2}(\mathbf{Z})$. Now, by Lemma 4.1.7, we know that this group has (Wohlfahrt) level dividing $2Np$. In particular $f(p\tau)$ has cusp width dividing $2Np$ at each cusp, and hence $f(p\tau)\in R_{2Np}$.
@@
Suppose that $(p,2N)=1$ is prime. Suppose that $f(\tau)\in R_{2N}$ is not invariant under a congruence subgroup. Then the form $f(p\tau)$ is not in the $M_{2Np}$-algebra generated by $R_{2N}$.
@@
Let $p$ be a prime not dividing $2N$. Suppose that $R_{2N}:M_{2N}]>1$. Then one has
$R_{2Np}:M_{2Np}]\geq 2[R_{2N}:M_{2N}].$
@@
Let $f(\tau)$ be a form in $R_{2Np}$ which is not in $M_{2N}$. By Theorem 4.3.1, we deduce that $f(p\tau)\in R_{2Np}$ is not in the $M_{2Np}$-algebra $R_{2N}M_{2Np}$ generated by $R_{2N}$ (which is a subfield of $R_{2Np}$). We have
$=$ $\ [R_{2Np}:R_{2N}M_{2Np}][R_{2N}M_{2Np}:M_{2N}]$
$=$ $\ [R_{2Np}:R_{2N}M_{2Np}][R_{2N}:M_{2N}][M_{2Np}:M_{2N}],$
because the intersection of $M_{2Np}$ and $R_{2N}$ is $M_{2N}$. Thus
$\frac{[R_{2Np}:M_{2Np}]}{[R_{2N}:M_{2N}]}=[R_{2Np}:R_{2N}M_{2Np}]$
@@
Our goal is to prove that $R_{2N}=M_{2N}$. As noted in Lemma 4.2.2, $R_{2N}:M_{2}]<\infty$. The degree of $M_{2N}$ over $M_{2}$ is equal to the degree of the modular curve $Y(2N)$ over $Y(2)$, and this is given, for $N>1$, by the explicit formula
(4.3.3) $=\frac{1}{2}[\Gamma(2):\Gamma(2N)]=$ $\ \frac{(2N)^{3}}{2[\mathrm{SL}_{2}(\mathbf{Z}):\Gamma(2)]}\prod_{p|2N}\left(1-\frac{1}{p^{2}}\right)$
$>$ $\ \frac{(2N)^{3}}{2[\mathrm{SL}_{2}(\mathbf{Z}):\Gamma(2)]}\prod_{p}\left(1-\frac{1}{p^{2}}\right)=\frac{2N^{3}}{3\zeta(2)}=\frac{4N^{3}}{\pi^{2}}.$
The factor of $1/2$ comes from the fact that $E=-I\in\Gamma(2)$ and the degree of $Y(2N)$ over $Y(2)$ is the index of the images of these groups inside $\mathrm{PSL}_{2}(\mathbf{Z})$. Since $R_{2N}:M_{2}]=[R_{2N}:M_{2N}]\cdot[M_{2N}:M_{2}]$, it follows that we have a bound:
(4.3.4) $R_{2N}:M_{2N}]\leq\frac{\pi^{2}[R_{2N}:M_{2}]}{4N^{3}}$
for all $N$. We can now compare this bound against the one coming from Theorem 4.3.2.
@@
$R_{2N}:M_{2}]\leq CN^{3}\log N$
for all integers $N$. Then $R_{2N}=M_{2N}$ for every $N$, that is, the unbounded denominators conjecture holds.
@@
Assume there exists an $N$ such that $R_{2N}\neq M_{2N}$. Let $S$ denote the set of primes <X<X which are co-prime to $2N$. By induction, Theorem 4.3.2 implies for such an $N$ that, for any $\varepsilon>0$,
(4.3.6) $R_{2N\prod_{p\in S}p}:M_{2N\prod_{p\in S}p}]\geq 2^{\#S}>2^{(1-\varepsilon)X/\log X},$
for sufficiently large $X$ (depending on $N$ and $\varepsilon$) by the prime number theorem. The right-hand side certainly increases faster than any power of $X$. On the other hand, from the assumed bound on $R_{2N}:M_{2}]$ together with the bound (4.3.4), we obtain
(4.3.7) $[R_{2N\prod_{p\in S}p}:M_{2N\prod_{p\in S}p}]\leq$ $\ \frac{C\pi^{2}}{4}\cdot\log\left(2N\prod_{p\in S}p\right)$
$=$ $\ \frac{C\pi^{2}}{4}\cdot\log 2N+\frac{C\pi^{2}}{4}\sum_{p\in S}\log p<\frac{C\pi^{2}}{4}\cdot X(1+\varepsilon),$
where the last inequality follows (with the same $\varepsilon>0$) once more from the prime number theorem for sufficiently large $X$. Combining the bounds (4.3.6) and (4.3.7) gives, for all sufficiently large $X$,
$2^{(1-\varepsilon)X/\log X}<\frac{C\pi^{2}}{4}\cdot X(1+\varepsilon),$
which (by some margin!) is a contradiction for any fixed $\varepsilon<1$. ∎
@@
The argument still works with a bound weaker than $R_{2N}:M_{2}]\ll N^{3}\log N$, although $R_{2N}:M_{2}]\ll N^{3+\varepsilon}$ would not be strong enough.
@@ 4.4. Amalgams and a non-abelian version of Ihara’s Lemma
In § 4.3, we introduced a group $G=G_{2N}\subset\langle E,\Gamma(2N)\rangle$ which was normal with finite index and had $L(G_{2N})=2N$. In this section, we consider more generally (up to a notational shift) a group $G=G_{N}\subset\langle E,\Gamma(N)\rangle$ which is normal of finite index and with $L(G_{N})=N$, and then apply our results to the particular group $G$ of § 4.3 when we prove of Theorem 4.3.1. (See equations (4.4.9) and (4.4.10) and the surrounding discussion.)
@@
This is more or less precisely [Tho89, Theorem 3] with the addition of level structure as in [Ber94]. Ihara’s Lemma [Rib84] is (informally) the statement that the two maps $H^{1}(\Gamma,\mathbf{F}_{q})\rightarrow H^{1}(\Gamma_{0}(p),\mathbf{F}_{q})$ coming from the restriction map and (respectively) the restriction map conjugated by $A$ have images which are as disjoint as possible. One may think of Lemma 4.4.1 as a non-abelian version of Ihara’s Lemma, because (as explained below in the proof of Lemma 4.6.2) the case when $S$ is a vector space over $\mathbf{F}_{q}$ reduces precisely to the statement of Ihara’s Lemma as proved by Ribet [Rib84]. (The proofs of both claims are very similar.)
@@
The intersection of $\langle E,\Gamma(N)\rangle$ with $A\langle E,\Gamma(N)\rangle A^{-1}$ is the group $\langle E,\Gamma(N)\rangle\cap\Gamma_{0}(p)$. We proceed by contradiction. Assume that the map $(f_{1},f_{2})$ is not surjective. By Goursat’s lemma, there exists a nontrivial quotient $\Delta$ of $S$ and projections $\pi_{i}:S\rightarrow\Delta$ such that the composites $\pi_{1}\circ f_{1}$ and $\pi_{2}\circ f_{2}$ agree. We define a map $g_{1}$ by the composite
@@
Let $\gamma=\left(\begin{matrix}1&N\\ 0&1\end{matrix}\right)$ and $\eta=\left(\begin{matrix}1&0\\ N&1\end{matrix}\right)$. The assumption that $L(G)=N$ and $G$ has finite index in $\mathrm{SL}_{2}(\mathbf{Z})$ impies that $\gamma,\eta\in G$. Since $A^{-1}\gamma^{p}A=\gamma\in A^{-1}GA\cap\Gamma_{0}(p)$, we see that $\gamma\in\ker(f_{1})$ and $\gamma\in\ker(f_{2})$, and yet
@@
(4.4.5) $H^{1}(\Gamma(N)/\mathrm{S}G,\mathbf{F}_{q})^{2}=H^{1}(S,\mathbf{F}_{q})^{2}\rightarrow H^{1}(\Gamma(N)\cap\Gamma_{0}(p),\mathbf{F}_{q}).$
@@
where $H^{1,\mathrm{cong}}(\Gamma(N),\mathbf{F}_{q})\subset H^{1}(\Gamma(N),\mathbf{F}_{q})$ denotes the classes which vanish after restriction to a congruence subgroup (Definition 4.5.1). This is because the kernel of any nontrivial map in $H^{1,\mathrm{cong}}(\Gamma(N),\mathbf{F}_{q})$ has level strictly divisible by $N$. The claim (4.4.5) follows from (4.4.6) as a consequence of Ihara’s Lemma, as proved by Ribet [Rib84] (see Lemma 4.6.2). The maps
$\langle E,\Gamma(N)\rangle\cap\Gamma_{0}(p)\rightarrow B/\langle U\rangle\rightarrow\mathbf{F}_{q}$
on the other hand come from the classes in $H^{1}(\Gamma(N)\cap\Gamma_{0}(p),\mathbf{F}_{q})$ which restricts to zero on $H^{1}(\Gamma(N)\cap\Gamma_{1}(p),\mathbf{F}_{q})$, and thus what is required is to upgrade the injection of (4.4.5) to an injection
(4.4.7) $H^{1}(\Gamma(N)/\mathrm{S}G,\mathbf{F}_{q})^{2}=H^{1}(S,\mathbf{F}_{q})^{2}\rightarrow H^{1}(\Gamma(N)\cap\Gamma_{1}(p),\mathbf{F}_{q}),$
@@
Because $\gamma$ and $E$ map to zero in $S\times S$ and $B/\langle E,U\rangle$ has order (p1)/2(p-1)/2, the proof of Lemma 4.4.4 is almost immediate if one imposes the additional hypothesis that $\left(\frac{p-1}{2},|S|\right)=1$. In particular, one would not have to appeal to the results in § 4.5 and § 4.6 (which are not used elsewhere in this paper). It turns out that proving Lemma 4.4.4 under this weaker hypothesis would suffice for the proof of the unbounded denominators conjecture. The key point is that if $\langle E,\Gamma(N)\rangle/G_{N}\simeq S$ and $G_{Np}$ is the group given by the intersection of the three groups $\langle E,\Gamma(Np)\rangle$, $G$, and $AGA^{-1}$, then $\langle E,\Gamma(Np)\rangle/G_{Np}\simeq S\times S$. In particular, one can control the primes dividing $S$ as one varies $N$. Then, in the argument of Proposition 4.3.5, instead of adding all primes <X<X prime to $N$, one only includes primes in some arithmetic progression satisfying the congruence $\left(\frac{p-1}{2},|S|\right)=1$ for some fixed $S$. However, it seems more natural to prove Lemma 4.4.4 without such an ugly hypothesis. Additionally, § 4.5 and § 4.6 may be of independent interest.
Returning to the assumptions of Lemma 4.4.4, let $K=\langle\ker((f_{1},\pi)),\ker(f_{2})\rangle$ be the group generated by $\ker((f_{1},\pi))$ and $\ker(f_{2})$. We deduce from Lemma 4.4.4 that the image of $(f_{1},f_{2},\pi)$ contains the elements $(x,0,z)$ and $(0,y,0)$ for any triple $(x,y,z)\in S\times S\times B$. But the pre-images of these elements clearly lie in $\ker(f_{2})$ and $\ker((f_{1},\pi))$ respectively, and thus lie in $K$. But then the pre-image of any element lies in $K$, and we deduce that $K=\langle E,\Gamma(N)\rangle\cap\Gamma_{0}(p)$, or equivalently that
@@
Now specializing to the group $G=G_{2N}$ of § 4.3, we obtain the corresponding identity
(4.4.10) $\langle E,G\cap\Gamma(2Np),A^{-1}GA\cap\Gamma_{0}(p)\rangle=\langle E,\Gamma(2N)\rangle\cap\Gamma_{0}(p).$
@@
Consider the function $f(p\tau)$. Assume that this lies in the algebra generated by $f(\tau)$ and $M_{2Np}$. Then $f(p\tau)$ is invariant under both $A^{-1}GA\cap\mathrm{SL}_{2}(\mathbf{Z})$ and $G\cap\Gamma(2Np)$. But from (4.4.10) we see that these groups together generate a congruence subgroup, and thus $f(p\tau)$ and $f(\tau)$ are congruence, a contradiction. ∎